Could someone please show me how to solve by elimination method?
x + 3y =23
-x +8y = 32

Found 3 solutions by rwm, richard1234, ewatrrr:
Answer by rwm(914)

add the two equations
11y=55
y=5
now sub 5 in for y
x+15=23
x=8

Answer by richard1234(7193)

Add the two equations, thereby "eliminating" x:
11y = 55
y = 5

Then, replace y = 5 into either equation to obtain x = 8.

Answer by ewatrrr(24785)


Hi
x + 3y = 23
-x + 8y = 32 |'Adding the two EQs' to eliminate x
11y = 55
y = 5 and x = 8 (23-3*5)